OpenStudy (anonymous):

Find the derivative of f(x) = 5 divided by x at x = -1.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (decentnabeel):

f(x)=5/x x=-1 =5/-1 =-5 so that the derivate of the constant is zero

OpenStudy (anonymous):

?? @DecentNabeel take derivative first, then replace x =-1 in

OpenStudy (anonymous):

f(x) =5 divided by x, means \(f(x) =\dfrac{5}{x}\)

OpenStudy (decentnabeel):

\[\frac{d}{dx}\left(\frac{5}{x}\right)=-\frac{5}{x^2}\]

OpenStudy (anonymous):

So the derivative is 5?

OpenStudy (anonymous):

or -5?

OpenStudy (decentnabeel):

hen x replace for -1 -5/(-1)^2 =-5/1 =-5

OpenStudy (anonymous):

ok